Please answer this: 963, 899, 843, 795, ?

Found 2 solutions by ikleyn, TigerAlgebra:
Answer by ikleyn(52788)
Answer by TigerAlgebra(40)

963-899=64
899-843=56
843-795=48
As the differences go down by 8 we can tll the next one will be 40
795-?=40
the number is 755 (I think)